# Phi-3 Safety Post-Training: Aligning Language Models with a “Break-Fix” Cycle

### v1

@@ Abstract
Recent innovations in language model training have demonstrated that it is possible to create highly performant models that are small enough to run on a smartphone. As these models are deployed in an increasing number of domains, it is critical to ensure that they are aligned with human preferences and safety considerations. In this report, we present our methodology for safety aligning the Phi-3 series of language models. We utilized a “break-fix” cycle, performing multiple rounds of dataset curation, safety post-training, benchmarking, red teaming, and vulnerability identification to cover a variety of harm areas in both single and multi-turn scenarios. Our results indicate that this approach iteratively improved the performance of the Phi-3 models across a wide range of responsible AI benchmarks.
@@ 1 Introduction
Given the computational cost associated with large language models (LLMs), there is increasing interest in developing smaller models with lower compute and memory requirements. Recent research has demonstrated that it is possible to create performant small language models (SLMs) by training on highly curated and synthetic datasets [GZA+23, LBE+23, JBA+23]. Microsoft recently released the Phi-3 series of SLMs, including Phi-3-mini (3.8B), Phi-3-small (7B), and Phi-3-medium (14B). For example, Phi-3-mini is small enough to run on a smartphone but achieves 69% on MMLU and 8.38 on MT-Bench, making it competitive with Mixtral 8x7B and GPT-3.5. [AJA+24].
@@ 2.3 Red Teaming
The Phi-3 release candidates were shared with a centralized Microsoft AI Red Team (AIRT), which leveraged a variety of techniques to test the models for responsible AI (RAI) risks across a range of categories. More specifically, the AIRT probed Phi-3 models for harmful content using both single-turn prompts and multi-turn conversations. These strategies were further split into “low-skilled adversary” and “intermediate adversary” personas, as summarized in Table 1. These personas were scoped based on the most common ways in which users might elicit harmful content from models. A low-skilled adversary was defined as a typical chatbot user who tries to generate harmful content by asking for it directly, while an intermediate adversary actively attempts to subvert a model’s safety guardrails using, for example, basic encodings and jailbreaks. To gauge the risk posed by Phi-3 models in comparison to open-source equivalents, the AIRT performed the same testing on Gemma-7B [TMH+24], Mixtral-8x7B [JSR+24], and Llama-3-In [TLI+23].
@@
The AIRT operated independently of the safety post-training team and probed for harmful content across a range of categories, including content related to current events, phishing and cybersecurity, fairness and bias, hate speech, sexual content, and violence. Note that these harm categories were selected in accordance with Microsoft’s Responsible AI Standard [Mic22] and were not necessarily the same as those covered by the preference datasets used for safety post-training.
@@
To red team the Phi-3 models at scale and cover as much of the risk surface as possible, the AIRT utilized PyRIT: Python Risk Identification Toolkit [LML+24]. [^1] PyRIT GitHub repository: https://github.com/Azure/PyRIT PyRIT is an open-source project that provides automation to support prompt generation, prompt conversion (e.g., to apply encodings and jailbreaks), response scoring, and even multi-turn conversations driven by an attacker LLM. To verify the accuracy of PyRIT automation, the AIRT manually checked the scored results and made corrections where necessary. Note that PyRIT was developed specifically to support red teaming operations and is separate from automation used to calculate responsible AI benchmarks. For the intermediate adversary multi-turn scenario, AIRT manually applied strategies like Crescendo, which uses seemingly benign prompts and gradually escalates the conversation to jailbreak a model [RSE24].
@@
In the next section, we present the results of the final Phi-3 models on a range of responsible AI benchmarks, along with the results achieved by comparable open-source models. We utilized these benchmarks to track overall safety performance and compare release candidates throughout the safety post-training process. It is important to note, however, that a model’s full risk profile can never be fully captured by a single set of metrics. In contrast with safety benchmarks, red teaming targets emerging harm areas, leverages the latest adversarial techniques, and can address ambiguous scenarios in which a model’s behavior might be interpreted in multiple ways. Importantly, red teaming centers the human elements of AI safety and can help answer questions related to how users might feel while interacting with a model. For example, in what scenarios might the model make users feel uncomfortable? Is the model at risk of providing dangerous or harmful advice? Are users likely to trust the model? During the red teaming phase of the break-fix cycle, questions like these played an important role in deciding where to focus further safety post-training efforts.
@@ 3.1 RAI Safety Benchmarks
Throughout the safety alignment process, we used a range of responsible AI (RAI) benchmarks – including both public datasets and Microsoft internal measurements – to track the safety performance of Phi-3 models and compare release candidates. In this section, we explain these benchmarks in detail and present the results achieved by the final release candidates, as well as those achieved by Mistral-7B, Gemma-7B, and Llama-3-In, for comparison.
@@ 4.1 Responsible Downstream Development
  • Quality of Service: The Phi-3 models are trained primarily on English text. Languages other than English will experience worse performance. English language varieties with less representation in the training data might experience worse performance than standard American English.
@@ 5 Conclusion
GenAI Model Team: Emman Haider, Daniel Perez-Becker, Thomas Portet, Piyush Madan, Amit Garg, David Majercak, Wen Wen, Dongwoo Kim, Ziyi Yang, Jianwen Zhang, Hiteshi Sharma

### v2

@@ Abstract
Recent innovations in language model training have demonstrated that it is possible to create highly performant models that are small enough to run on a smartphone. As these models are deployed in an increasing number of domains, it is critical to ensure that they are aligned with human preferences and safety considerations. In this report, we present our methodology for safety aligning the Phi-3 series of language models. We utilized a “break-fix” cycle, performing multiple rounds of dataset curation, safety post-training, benchmarking, red teaming, and vulnerability identification to cover a variety of harm areas in both single and multi-turn scenarios. Our results indicate that this approach iteratively improved the performance of the Phi-3 models across a wide range of responsible AI benchmarks. Finally, we include additional red teaming strategies and evaluations that were used to test the safety behavior of Phi-3.5-mini and Phi-3.5-MoE, which were optimized for multilingual capabilities.
@@ 1 Introduction
Given the computational cost associated with large language models (LLMs), there is increasing interest in developing smaller models with lower compute and memory requirements. Recent research has demonstrated that it is possible to create performant small language models (SLMs) by training on highly curated and synthetic datasets [GZA+23, LBE+23, JBA+23]. In April 2024, Microsoft released the Phi-3 series of SLMs, including Phi-3-mini (3.8B), Phi-3-small (7B), and Phi-3-medium (14B). For example, Phi-3-mini is small enough to run on a smartphone but achieves 69% on MMLU and 8.38 on MT-Bench, making it competitive with Mixtral 8x7B and GPT-3.5. [AJA+24].
@@ 2.3 Red Teaming
2.3.1 English-language testing (Phi-3)
The Phi-3 release candidates were shared with a centralized Microsoft AI Red Team (AIRT), which leveraged a variety of techniques to test the models for responsible AI (RAI) risks across a range of categories. Because the original Phi-3 models were not trained on curated multilingual datasets, this red teaming was conducted in English.
AIRT probed Phi-3 models for harmful content using both single-turn prompts and multi-turn conversations. These strategies were further split into “low-skilled adversary” and “intermediate adversary” personas, as summarized in Table 1. Personas were scoped based on the most common ways in which users might elicit harmful content from models. A low-skilled adversary was defined as a typical chatbot user who tries to generate harmful content by asking for it directly, while an intermediate adversary actively attempts to subvert a model’s safety guardrails using, for example, basic encodings and jailbreaks. To gauge the risk posed by Phi-3 models in comparison to open-source equivalents, AIRT performed the same testing on Gemma-7B [TMH+24], Mixtral-8x7B [JSR+24], and Llama-3-In [TLI+23].
@@ 2.3.1 English-language testing (Phi-3)
AIRT operated independently of the safety post-training team and probed for harmful content across a range of categories, including content related to current events, phishing and cybersecurity, fairness and bias, hate speech, sexual content, and violence. Note that these harm categories were selected in accordance with Microsoft’s Responsible AI Standard [Mic22] and were not necessarily the same as those covered by the preference datasets used for safety post-training.
@@
To red team the Phi-3 models at scale and cover as much of the risk surface as possible, AIRT utilized PyRIT: Python Risk Identification Toolkit [^1] PyRIT GitHub repository: https://github.com/Azure/PyRIT [LML+24]. PyRIT is an open-source project that provides automation to support prompt generation, prompt conversion (e.g., to apply encodings and jailbreaks), response scoring, and even multi-turn conversations driven by an attacker LLM. To verify the accuracy of PyRIT automation, AIRT manually checked the scored results and made corrections where necessary. Note that PyRIT was developed specifically to support red teaming operations and is separate from automation used to calculate responsible AI benchmarks. For the intermediate adversary multi-turn scenario, AIRT manually applied strategies like Crescendo, which uses seemingly benign prompts and gradually escalates the conversation to jailbreak a model [RSE24].
@@
In section 3.1, we present the results of the final Phi-3 models on a range of responsible AI benchmarks, along with the results achieved by comparable open-source models. We utilized these benchmarks to track overall safety performance and compare release candidates throughout the safety post-training process. It is important to note, however, that a model’s full risk profile can never be fully captured by a single set of metrics. In contrast with safety benchmarks, red teaming targets emerging harm areas, leverages the latest adversarial techniques, and can address ambiguous scenarios in which a model’s behavior might be interpreted in multiple ways. Importantly, red teaming centers the human elements of AI safety and can help answer questions related to how users might feel while interacting with a model. For example, in what scenarios might the model make users feel uncomfortable? Is the model at risk of providing dangerous or harmful advice? Are users likely to trust the model? During the red teaming phase of the break-fix cycle, questions like these played an important role in deciding where to focus further safety post-training efforts.
2.3.2 Multilingual testing (Phi-3.5)
In August 2024, Microsoft announced Phi-3.5-mini [^2] Phi-3.5-mini model card: https://huggingface.co/microsoft/Phi-3.5-mini-instruct and Phi-3.5-MoE. [^3] Phi-3.5-MoE model card: https://huggingface.co/microsoft/Phi-3.5-MoE-instruct In addition to performance and latency improvements, these models provide multilingual capabilities, and AIRT conducted an additional operation to probe Phi-3.5-MoE for violent content, sexual content, and hate speech across four languages: Chinese, Spanish, Dutch, and English (baseline).
Given the limited precedent for multilingual RAI red teaming, AIRT enlisted fluent speakers in these languages to perform red teaming manually. In particular, the following techniques were leveraged:
  • Single-turn prompting: To compare the behavior of Phi-3.5 in English, Chinese, Dutch, and Spanish, AIRT manually translated and evaluated the model with the same set of prompts across four harm categories – violent content, sexual content, ECI, and hate speech.
  • Single-turn jailbreaks: To compare the robustness of Phi-3.5 to jailbreaks in English, Chinese, Dutch, and Spanish, AIRT leveraged a dataset of well-known jailbreaks (e.g., BetterDAN, AntiGPT, etc.) and applied these jailbreaks to a variety of prompts that the model typically refuses.
  • Multiturn role-playing: To compare the robustness of Phi-3.5 in English, Chinese, Dutch, and Spanish to multiturn strategies, AIRT manually applied a role-playing strategy in both generic and culturally specific contexts.
Overall, AIRT found that the safety behavior of Phi-3.5-MoE is similar in the four languages tested. More specifically, the model typically refused direct requests for harmful content and was robust to common jailbreaks in all languages. Although slightly higher refusal rates were observed in English than other languages, there was no significant difference in terms of harmful content generated. Note, however, that the model often provided English refusals even when the request for harmful content was in another language.
AIRT found that Phi-3.5-MoE model was susceptible to a multiturn role-playing strategy, which was used to elicit sexual and violent content, fake news, and hate speech in all languages, including English. This multiturn strategy was also applied to culturally specific scenarios in Dutch and Chinese, but the model generated similar content when prompted in English.
Note that AIRT did not perform extensive testing of medium and low-resource languages, and it is possible that safety post-training in English does not transfer as effectively to other languages and scenarios.
@@ 3.1 RAI Safety Benchmarks
Throughout the safety alignment process, we used a range of responsible AI (RAI) benchmarks, including both public datasets and Microsoft internal measurements, to track the safety performance of Phi-3 models and compare release candidates. In this section, we explain these benchmarks in detail and present the results achieved by the final release candidates, as well as those achieved by Mistral-7B, Gemma-7B, and Llama-3-In, for comparison.
@@ 3.2 Iterative Safety Alignment
3.3 Multilingual RAI Safety Benchmarks
A multi-faceted approach was adopted to evaluate the Phi-3.5 language models for safety behaviors, with additional measures taken to account for the multilingual capabilities introduced by this release. Our approach to multilingual safety evaluation included both internal automated benchmarks and open-source benchmarks. In this section, we explain these benchmarks and present the results achieved by the final release candidates as well as those achieved by Mistral-Nemo.
3.3.1 Internal Automated Measurement
First, we reproduced the English-language measurements described in section 3.1.1 for the Phi-3.5 models. The results can be seen in Table 6. We found, for the vast majority of metrics, that our 3.5 releases (Phi-3.5-mini-instruct and Phi-3.5-MoE-instruct) were on par with or better than the baseline model (Mistral-Nemo).
Phi-3.5-mini-instruct 3.8b Phi-3.5-MoE-instruct 16$\times$ 3.8b Mistral-Nemo 12b Ungroundedness 0.459 0.271 0.244 Third Party Harm (DR-1) 0.243 0.152 0.599 Harmful Content Continuation (DR-3) 0.006 0.004 0.025 Harmful Content Summarization (DR-3) 0.119 0.120 0.110 Jailbreak (DR-1) 0.119 0.110 0.174
Table 6: Results of Phi-3.5 models and baseline model on the Microsoft internal multi-turn conversation benchmarks (English-only). Note that a lower value indicates a better performance for all metrics in the table.
To evaluate the risk of multilingual harmful content generation, we augmented our existing internal benchmark for additional languages supported by the Phi-3.5 models. Table 7 shows the harmful content continuation metric for some of the languages supported by Phi-3.5. We observe better performance than Mistral-Nemo, with Phi-3.5-MoE-instruct tending to do slightly better than Phi-3.5-mini-instruct.
We are continuing research and investments into increasing our multilingual benchmark coverage for additional metrics and languages.
Harmful Content Continuation (DR-3) Phi-3.5-mini-instruct 3.8b Phi-3.5-MoE-instruct 16$\times$ 3.8b Mistral-Nemo 12b Chinese Simplified 0.015 0.007 0.026 French 0.010 0.010 0.034 German 0.007 0.010 0.027 Italian 0.009 0.009 0.023 Japanese 0.010 0.008 0.024 Portuguese (Brazil) 0.012 0.017 0.026 Spanish 0.016 0.008 0.026
Table 7: Results of Phi-3.5 models and baseline model on the “Harmful Content Continuation” Microsoft internal multi-turn conversation benchmark across several languages. Because the reported metric is DR-3 (explained in section 3.1.1), a lower value indicates better performance.
3.3.2 XSafety
To measure refusal rates, we used the XSafety dataset and methodology from [WTC+24] since the XSTest dataset described in section 3.1.2 was not multilingual. The XSafety dataset consists of $\sim$2,800 malicious prompts per language, and we use GPT-4 to determine whether Phi-3.5’s responses to these prompts constitute refusals or not.
Results are given in Table 8 and show that Phi-3.5-mini-instruct and Phi-3.5-MoE-instruct perform almost as well or even better than Mistral-Nemo.
Refusal Rate Phi-3.5-mini-instruct 3.8b Phi-3.5-MoE-instruct 16$\times$ 3.8b Mistral-Nemo 12b Chinese Simplified 0.691 0.659 0.706 French 0.874 0.865 0.870 German 0.881 0.880 0.865 Japanese 0.600 0.714 0.673 Spanish 0.864 0.878 0.853
Table 8: Results of Phi-3.5 models and baseline model on the XSafety benchmark across various languages. A higher value indicates better performance.
3.3.3 RTP-LX
To assess the ability of the model to detect toxic speech, we used the framework described in [dWWA+24]. The dataset consists of $\sim$1,000 prompts per language, along with a toxicity level provided by human reviewers. We then asked Phi-3.5 to rate the toxicity of the prompts and compared the Phi-3.5 ratings to human ratings. This is similar to what we did in section 3.1.4 for English with the Toxigen dataset.
Results can be found in Table 9 and indicate that Phi-3.5-MoE-instruct can detect toxic speech almost as well as Mistral-Nemo, and better than Phi-3.5-mini-instruct.
Kappa/Inter-annotator Agreement Phi-3.5-mini-instruct 3.8b Phi-3.5-MoE-instruct 16$\times$ 3.8b Mistral-Nemo 12b Chinese Simplified 0.195 0.287 0.323 French 0.211 0.377 0.349 German 0.162 0.251 0.276 Italian 0.172 0.295 0.323 Japanese 0.206 0.285 0.315 Portuguese (Brazil) 0.203 0.323 0.366 Spanish 0.196 0.278 0.297
Table 9: Results of Phi-3.5 models and baseline model on the RTP-LX benchmark across various languages. A higher value indicates better performance.
@@ 4.1 Responsible Downstream Development
  • Quality of Service: The Phi-3 models are trained primarily on English text. Languages other than English will experience worse performance. English language varieties with less representation in the training data might experience lower performance than standard American English.
  • Multilingual Performance & Safety Gaps: We believe it is important to make language models more widely available across different languages, but the Phi-3 models still exhibit challenges common across multilingual releases. As with any deployment of LLMs, developers will be better positioned to test for performance or safety gaps for their linguistic and cultural context and customize the model with additional fine-tuning and appropriate safeguards.
@@
  • Long Conversation: Phi-3 models, like other models, can in some cases generate responses that are repetitive, unhelpful, or inconsistent in very long chat sessions in both English and non-English languages. Developers are encouraged to place appropriate mitigations, like limiting conversation turns to account for possible conversational drift.
@@ 5 Conclusion
GenAI Model Team: Emman Haider, Daniel Perez-Becker, Thomas Portet, Piyush Madan, Amit Garg, Atabak Ashfaq, David Majercak, Wen Wen, Dongwoo Kim, Ziyi Yang, Jianwen Zhang, Hiteshi Sharma